# SpreadsheetLLM: Encoding Spreadsheets for Large Language Models

### v1

@@ 6 Conclusion
While our SpreadsheetLLM frameworks have markedly advanced how LLMs interpret and utilize spreadsheets, they also illuminate areas ripe for further research and development. Currently, our methods do not yet harness spreadsheet format details such as background color and borders, because they take too many tokens. However, these elements often contain valuable contextual and visual cues that could further refine our understanding and processing of spreadsheet data. Additionally, while SheetCompressor effectively aggregates data regions, it does not currently employ a sophisticated semantic-based compression method for cells containing natural language. For example, categorizing terms like "China," "America," and "France" under a unified label such as "Country" could not only increase the compression ratio but also deepen the semantic understanding of the data by LLMs. Exploring these advanced semantic compression techniques will be a key focus of our ongoing efforts to enhance the capabilities of SpreadsheetLLM.
@@ Appendix B Traditional Encoding Methods for Spreadsheets
To quantitatively assess these methods, we conducted ICL experiments using the GPT-4 model on spreadsheet detection tasks. The results, detailed in Table 6, confirmed that while the Markdown method outperformed XML and HTML in terms of lower token usage, it still fell short in addressing the needs of spreadsheet encoding effectively.
@@ Appendix C Lightweight Heuristics for Structural-anchor Proposal
Finally, we take the remaining candidate boundaries to derive structural anchors. However, due to the challenge of fine-grained discriminating headers, titles, and notes for heuristics, the candidate boundaries produced by the above heuristics only achieve 46.3% F1 score in EoB-0 metric and 65.0% EoB-2 metric in our boundary detection test set. Fortunately, including neighboring rows and columns largely increases the coverage of real bounding rows and columns, because headers, titles, and notes usually span few rows and columns. So we propose to not only use the exact bounding rows and columns as structural anchors but also include rows and columns within $k$ rows and columns neighboring the structural anchors to preserve titles, notes, and headers as much as possible. This allows LLMs to further determine the exact boundaries by leveraging their advanced capabilities in semantic understanding and reasoning. When $k$ is set to 4, over 97% of the border rows and columns in ground truth tables are preserved. This ensures that structure anchors rarely lose critical information of the table skeleton. The source code and an executable tool to use this method can be found in the supplements to improve reproducibility.
@@ Appendix E Spreadsheet Table Detection Test Dataset Quality Improvement Pipeline
As a result, we obtained a well-annotated dataset with 167 spreadsheets containing 268 tables for type 1, 21 spreadsheets with 43 tables for type 2, and 10 spreadsheets for type 3. All the raw files and labels of the test set are attached to the supplements. We selected data labeled as type 1 and type 2 as the test set, comprising a total of 188 entries.
@@ Appendix G Experiment Setup
We have attached the fine-tuned file and parameters in the Supplementary materials.
@@ Appendix I Cost calculation
We use the ICL price of GPT4 due to the absence of fine-tuned GPT4’s price. We neglect the output sequence since it is much shorter than the input sequence in tasks like spreadsheet boundary detection and QA. The average cost of processing a spreadsheet in our test set has decreased to \$0.000157 (62000/198 * 0.0005 / 1000) from \$0.00391 (1548000/198 * 0.0005 / 1000) for the GPT3.5 turbo, and to \$0.00939 (62000/198 * 0.03 / 1000) from \$0.235 (1548000/198 * 0.03 / 1000) for the GPT4, saving an impressive 96.0% in costs. The cost reduction similarly applies to all LLMs we used.
@@
Table 10: Compression performance on train & valid Datasets.
@@
The corresponding algorithm steps is shown in Algorithm 1.

### v2

@@ 6 Conclusion
While SpreadsheetLLM has markedly advanced how LLMs interpret and utilize spreadsheets, they also illuminate areas ripe for further research and development. Currently, our methods do not yet harness spreadsheet format details such as background color and borders, because they take too many tokens. However, these elements often contain valuable contextual and visual cues that could further refine our understanding and processing of spreadsheet data. Additionally, while SheetCompressor effectively aggregates data regions, it does not currently employ a sophisticated semantic-based compression method for cells containing natural language. For example, categorizing terms like "China," "America," and "France" under a unified label such as "Country" could not only increase the compression ratio but also deepen the semantic understanding of the data by LLMs. Exploring these advanced semantic compression techniques will be a key focus of our ongoing efforts to enhance the capabilities of SpreadsheetLLM.
@@ Appendix B Traditional Encoding Methods for Spreadsheets
To quantitatively assess these methods, we conducted ICL experiments using the GPT4 model on spreadsheet detection tasks. The results, detailed in Table 6, confirmed that while the Markdown method outperformed XML and HTML in terms of lower token usage, it still fell short in addressing the needs of spreadsheet encoding effectively.
@@ Appendix C Lightweight Heuristics for Structural-anchor Proposal
Finally, we take the remaining candidate boundaries to derive structural anchors. However, due to the challenge of fine-grained discriminating headers, titles, and notes for heuristics, the candidate boundaries produced by the above heuristics only achieve 46.3% F1 score in EoB-0 metric and 65.0% EoB-2 metric in our boundary detection test set. Fortunately, including neighboring rows and columns largely increases the coverage of real bounding rows and columns, because headers, titles, and notes usually span few rows and columns. So we propose to not only use the exact bounding rows and columns as structural anchors but also include rows and columns within $k$ rows and columns neighboring the structural anchors to preserve titles, notes, and headers as much as possible. This allows LLMs to further determine the exact boundaries by leveraging their advanced capabilities in semantic understanding and reasoning. When $k$ is set to 4, over 97% of the border rows and columns in ground truth tables are preserved. This ensures that structure anchors rarely lose critical information about the table skeleton.
@@ Appendix E Spreadsheet Table Detection Test Dataset Quality Improvement Pipeline
As a result, we obtained a well-annotated dataset with 167 spreadsheets containing 268 tables for type 1, 21 spreadsheets with 43 tables for type 2, and 10 spreadsheets for type 3. We selected data labeled as type 1 and type 2 as the test set, comprising a total of 188 entries.
@@ Appendix G Experiment Setup
lora_dim=32,n_epochs=-1,batch_size=-1,learning_rate_multiplier=1,weight_decay_multiplier=1e-05,prompt_loss_weight=0,trim_mode=right.
@@ Appendix I Cost calculation
We use the ICL price of GPT4 due to the absence of fine-tuned GPT4’s price. We neglect the output sequence since it is much shorter than the input sequence in tasks like spreadsheet boundary detection and QA. The average cost of processing a spreadsheet in our test set has decreased by 96.0% for GPT3.5 turbo and GPT4, saving an impressive 96.0% in costs. The cost reduction similarly applies to all LLMs we used.
@@
Table 10: Compression performance on train & valid Datasets.
@@
The corresponding algorithm steps are shown in Algorithm 1.